OpenStudy (anonymous):

Find the solution of the IVP: y''+4y=t^2+3e^t, y(0)=0, y'(0)=2.

OpenStudy (zzr0ck3r):

can you find the characteristic equation, and the Eigen values?

OpenStudy (anonymous):

now you need to find the particular solution of the form At^2+Bt+C+De^t

OpenStudy (zzr0ck3r):

(at² + bt + c + de^t)'' + 4(at² + bt + c + de^t) = t² + 3e^t

OpenStudy (zzr0ck3r):

now can you get a,b,c,d?

OpenStudy (zzr0ck3r):

this is like partial decomp...

OpenStudy (zzr0ck3r):

@Cacastro ? do you see what I am doing? we solved the homogenous version first, now we make up for the forcing terms

OpenStudy (zzr0ck3r):

the general non hg solution is c_1 cos(2t) + c_2 sin(2t) + at² + bt + c + de^t then we use initial to find c_1 and c_2

OpenStudy (zzr0ck3r):

(at² + bt + c + de^t)'' + 4(at² + bt + c + de^t) = t² + 3e^t 4at² + 4bt + (2a + 4c) + 5de^t = t² + 3e^t

OpenStudy (anonymous):

I'm trying to find a,b, c and d

OpenStudy (anonymous):

do you see that he simplified everything up to this \[4at ^{2}+4bt+(2a+4c)+5de ^{t}=t ^{2}+3e ^{t}\] can you look and match the coefficients to find a and b and d?

OpenStudy (zzr0ck3r):

4a=1 4b=0 2a+4c=0 5d=3

OpenStudy (anonymous):

a=1/4 , b=0, c=-1/8, d=1

OpenStudy (anonymous):

think you need to check d

OpenStudy (anonymous):

5d=3

OpenStudy (anonymous):

d=3/5

OpenStudy (anonymous):

good, now plug those values into the formula above and you'll have the particular solution

OpenStudy (zzr0ck3r):

not yet...you need to solve for your other coefficiants.

OpenStudy (zzr0ck3r):

coefficients

OpenStudy (anonymous):

the particular solution is not the final answer

OpenStudy (anonymous):

the solution geral is \[y(t)= C _{1} \cos 2t + C _{2} sen 2t + \frac{ 1 }{ 4 } t^2- \frac{ 1 }{ 8 } + \frac{ 3 }{ 5 }e^t\]

OpenStudy (anonymous):

and now i have to replace for the initial conditions

OpenStudy (anonymous):

yes?

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

I found C1= -19/40 and c2= 7/310 A solução do PVI é \[- \frac{ 19 }{40 } \cos 2t + \frac{ 7 }{ 10 } \sin 2t + \frac{ 1 }{ 4 }t^2 - \frac{ 1 }{ 8 } + \frac{ 3 }{ 5 } e^t \]

OpenStudy (anonymous):

guess i better do it to check your work, give me a few mins ok?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I got something different

OpenStudy (anonymous):

oh wait i forgot the 1/8

OpenStudy (anonymous):

that looks right nice work

OpenStudy (zzr0ck3r):

hmm I got 28/40 on sin

OpenStudy (zzr0ck3r):

nm lol

OpenStudy (zzr0ck3r):

sorry

OpenStudy (zzr0ck3r):

you got what I got, nice work...these are long:)

OpenStudy (anonymous):

yeah these can get really really messy;

OpenStudy (anonymous):

Thank very much